OpenStudy (anonymous):

If θ is an acute angle of a right triangle and cot θ = 12/5, find csc θ.

OpenStudy (anonymous):

I got it. Wait

OpenStudy (anonymous):

\[\text{Csc}\left[\text{ArcCot}\left[\frac{5}{12}\right]\right]=1.08333 \]

OpenStudy (anonymous):

Not right. wait

OpenStudy (anonymous):

thanks for your answer

OpenStudy (anonymous):

csc = 13/5

OpenStudy (anonymous):

that right

OpenStudy (anonymous):

i got it

OpenStudy (anonymous):

My posting is incorrect. This is what happens when one tries to rush out an answer.

OpenStudy (anonymous):

Sometimes that happens to me too. You tried to help, and that is important!

OpenStudy (anonymous):

\[\text{Csc}\left[\text{ArcCot}\left[\frac{12}{5}\right]\right]=\frac{13}{5} \]which agrees with emunrradtvamg